Find a particular solution to y + 6y + 5y = 6te2t.
Answer #1

Y+64 +5y = 6 + 027 (02+60+5)y= 67e27 Solution: Given D.C. Ds d dt Pt is Jp = 6tpat 22460+5 T. 6 t tezt keht (D+1) (0+5) (D+2